OpenStudy (anonymous):

What is the first step that you have to do in order to solve the proportion: one-third equals the fraction y over 26? A. 8.5 B. 26y = 3 C. y = 3(26) D. 3(y) = 1(26)

OpenStudy (anonymous):

1/3 = y/26?

OpenStudy (anonymous):

(question mark is there in the equation)

OpenStudy (anonymous):

1/3 = y/26 3 * y = 26 * 1 3y = 26 y = 26/3 y = ??

OpenStudy (anonymous):

y = 8.7

hartnn (hartnn):

first step is to cross multiply. 3y = 26*1 which is D

OpenStudy (anonymous):

the first step is cross multiply

OpenStudy (anonymous):

i wish i could give you all medals ! thanks so much for the help <3

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

kkk ::))